OpenStudy (shelld93):

Solve for t, include all working. 3t - 6 = 2t + 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (shelld93):

i worked up to 3t -6 +6 = 2t + 6 + 6 3t = 2t + 12 3t/3 = 2t + 6/3 but from there if that is even right i have no idea

OpenStudy (skullpatrol):

The whole idea is to get the variable ALONE on one side of the equation, so make that your goal

OpenStudy (skullpatrol):

You want to end up with t=?

OpenStudy (shelld93):

yeah i can get the lhs to t but theres the t on the other side :/

OpenStudy (anonymous):

@ShellD93 I THINK YOU ARE AGIAN NOT PUTTING BRACES WHILE DIVIDING BY 3 BOTH SIDES.. 3t/3 = 2t + 6/3

OpenStudy (shelld93):

far out.

OpenStudy (skullpatrol):